An airplane is in level flight at a velocity of 300 miles per hour and an altitude of 5000 ft when it drops a bomb. (a) How long does the bomb take to reach the ground? (b) What will the bomb's velocity be on impact?
1
Expert's answer
2020-11-12T09:36:24-0500

The initial speed in horizontal direction is "v_{0h} = 300\space miles/h = 134.112\space m/s". This speed does not change with time. The initial speed in vertical direction is "v_{0v} = 0m\/s". Thus, the distance travelled in vertial direction is



"h = \dfrac{gt^2}{2}"

where if "h = 3000ft = 914.4m" then "t" is the time of falling. Thus:



"t = \sqrt{\dfrac{2h}{g}} = \sqrt{\dfrac{2\cdot 914.4}{9.81}} \approx 13.65s"

The speed of the bomb in the vertical direction after time "t" will be


"v_v = gt = 9.81\times 13.65\approx 133.9065\space m/s"

The total speed


"v = \sqrt{v_v^2 + v_{0h}^2} = \sqrt{(133.9065)^2 + (134.112)^2} \approx 189.5m/s"

Answer. "t = 13.65s", and "v = 189.5 m\/s".